6.

Find the values of a and b, where a and b are real, given that (a + bi)(2 – i) = 5 – i.


1
Expert's answer
2021-07-19T05:43:07-0400
"(a+bi)(2-i)=2a-ai+2bi+b"

"=(2a+b)+(-a+2b)i=5-i"

"\begin{matrix}
 2a+b=5\\
 -a+2b=-1
\end{matrix}"

"\begin{matrix}
 b=-2a+5\\
 -a-4a+10=-1
\end{matrix}"

"\begin{matrix}
 a=2.2\\
 b=0.6
\end{matrix}"